OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 2 -1 1 -2

OpenStudy (anonymous):

1) firstly, since theyre asking you for X, you can cancel out y by multiplying the first equation by 4 and the second one by 3 so that both the Y terms become 12 and can cancel. so... --> (2x + 3y = 1) X4 (3x - 4y = 10) X3 2) therefore you get: 8x + 12y = 4 + 9x - 12y = 30 _________________ 3) after adding the equations, * i added them because 12 + (-12) will cancel to give 0* you end up getting: ---> 17x = 34 4)divide both sides by 17 to get x...therefore 34/17 =2 5) so x = 2